A couple bought some stock for $30 per share that pays an annual dividend of $1.80 per share. After 1 year the price of the stock was $33. Find the simple interest rate on the growth of their investment.
Answer by jorel1380(3719)

The initial price of the stock is $30. Then the growth of their investment would be:
33/30=1.1
1.1-1=.1, or 10% growth on their stock.
If you count the dividend as growth, then you have:
33+1.8/30=1.16, or 16% growth. ☺☺☺☺